# PartSAM: A Scalable Promptable Part Segmentation Model Trained on Native 3D Data

### v1

@@ Abstract
Segmenting 3D objects into parts is a long-standing challenge in computer vision. To overcome taxonomy constraints and generalize to unseen 3D objects, recent works turn to open-world part segmentation. These approaches typically transfer supervision from 2D foundation models, such as SAM, by lifting multi-view masks into 3D. However, this indirect paradigm fails to capture intrinsic geometry, leading to surface-only understanding, uncontrolled decomposition, and limited generalization. We present PartSAM, the first promptable part segmentation model trained natively on large-scale 3D data. Following the design philosophy of SAM, PartSAM employs an encoder–decoder architecture in which a triplane-based dual-branch encoder produces spatially structured tokens for scalable part-aware representation learning. To enable large-scale supervision, we further introduce a model-in-the-loop annotation pipeline that curates over five million 3D shape–part pairs from online assets, providing diverse and fine-grained labels. This combination of scalable architecture and diverse 3D data yields emergent open-world capabilities: with a single prompt, PartSAM achieves highly accurate part identification, and in a “Segment-Every-Part” mode, it automatically decomposes shapes into both surface and internal structures. Extensive experiments show that PartSAM outperforms state-of-the-art methods by large margins across multiple benchmarks, marking a decisive step toward foundation models for 3D part understanding. Our code and model will be released soon.
@@ 1 Introduction
Second, we propose an effective encoder that can scale to large 3D datasets. Specifically, we encode shapes into a triplane-based feature field (Chan et al., 2022) using a duan-branch transformer (Vaswani et al., 2017). To better leverage existing priors, we leverage one learnable branch scale to large 3D datasets, while the other frozen path preserves 2D priors of SAM learned through contrastive learning (Liu et al., 2025). Additionally, by incorporating input attributes beyond just coordinates, the model enhances the representation of local shape details.
@@ 2.1 Closed-world 3D Part Segmentation
Early learning-based methods (Qi et al., 2017; Thomas et al., 2019; Wang et al., 2019b; Zhang & Wonka, 2021) typically formulate part segmentation as point-level semantic or instance classification on 3D shapes. These approaches are commonly trained on ShapeNet-Part (Yi et al., 2016) and PartNet (Mo et al., 2019). The limited diversity of these datasets in terms of object and part categories, however, constrains their generalization to unseen data.
@@ 3 Method
An overview of PartSAM is shown in Figure 3. Our goal is to develop a scalable framework for segmenting arbitrary 3D parts, analogous to SAM (Kirillov et al., 2023) in the image domain. The architecture consists of two main components: an input encoder that transforms 3D shapes into structured feature embeddings, and a prompt-guided mask decoder that predicts segmentation masks conditioned on user prompts. We primarily consider 3D click points as prompts, including both positive and negative ones. Formally, given a 3D shape represented as a point cloud $P_{in}\subseteq\mathbb{R}^{N\times d}$, where each point may include 3D coordinates, surface normals, and optional RGB color (i.e., $d=9$), and a set of user prompts $P_{prompt}\subseteq\mathbb{R}^{N_{p}\times 3}$, PartSAM first transform the input shape into feature embeddings $F_{c}$ through the input encoder. These are combined with prompt embeddings $F_{p}$ in the mask decoder to generate binary segmentation masks $M_{out}\subseteq\mathbb{R}^{N\times 1}$. To fully realize this design at scale, we further introduce a data curation pipeline that mines diverse and fine-grained 3D part annotations from large-scale assets, enabling PartSAM to generalize across open-world scenarios.
@@ 3.2 Prompt-Guided Mask Decoder
For mask decoding, two special tokens are introduced: an output token $T_{out}$ that generates segmentation masks, and an IoU token $T_{iou}$ that estimates mask quality. These tokens, together with the prompt embeddings $F_{p}$, are concatenated and attend bidirectionally to the patch embeddings $F_{c}$:
@@ 3.3 Training Strategy
Following SAM (Kirillov et al., 2023), during training, we simulate interactive segmentation by first sampling a prompt from the center of the ground-truth mask and then iteratively sampling subsequent prompts from prediction error regions for 9 iterations. The segmentation masks are supervised with a combination of focal loss (Lin et al., 2017) and dice loss (Milletari et al., 2016). For each ground-truth mask, the triplet contrastive loss in Liu et al. (2025) is also used to strengthen the representation ability of the encoder:
$\mathcal{L}=\mathcal{L}_{focal}(M_{out},M_{gt})+\mathcal{L}_{dice}(M_{out},M_{gt})+\lambda\,\mathcal{L}_{triplet},$ (3)
where $M_{gt}$ is the ground-truth mask and $\lambda$ balances the triplet loss contribution.
@@ 4.2 Comparison of Class-Agnostic Part Segmentation
As introduced in Section 3.4, PartSAM supports automatic shape decomposition through its “segment every part” mode. We compare PartSAM against five state-of-the-art methods (Liu et al., 2023; Ma et al., 2024; Yang et al., 2024b; Tang et al., 2024; Liu et al., 2025) on this task. We use the PartObjaverse-Tiny (Yang et al., 2024b) and PartNet-E (Liu et al., 2023) datasets for quantitative evaluation and additionally use AI-generated meshes produced by Hunyuan3D (Lai et al., 2025) for qualitative comparison. We follow the experimental protocol from Liu et al. (2025) with a key modification to prevent potential label leakage from mesh connectivity. Specifically, since artist-created meshes in PartObjaverse-Tiny contain strong connectivity cues that align with ground-truth part labels (see Section A.2.4), we disable this connectivity information when evaluating PartField (Liu et al., 2025) and instead apply K-Means clustering directly on mesh faces. In contrast, SAMesh (Tang et al., 2024), which fundamentally depends on mesh connectivity through community detection, is evaluated in its original setting. Notably, for baselines that produce multi-scale part segmentations, we evaluate performance by computing the IoU for all candidate masks and reporting the best score. For our PartSAM, multi-scale segmentations are obtained by varying the NMS threshold value $T$ at $0.1$, $0.3$, $0.5$, and $0.7$, respectively.
Qualitative comparisons in Figure 7 further reveal that PartSAM produces robust part segmentations for diverse 3D shapes. By contrast, SAMesh often generates over-segmented outputs, as its direct projection and merging of 2D SAM masks into 3D space fails to capture underlying geometric structures. SAMPart3D and PartField, which rely on feature clustering, encounter difficulties on complex and detailed structures. Their clustering process frequently yields fragmented or semantically meaningless parts, for example, incorrectly partitioning the body of a grenade into several disjoint segments. We also conduct a comparison with PartField in the context of hierarchical multi-scale segmentation on AI-generated meshes (Figure 8). PartSAM delivers finer-grained results with sharper boundaries, highlighting its capability to produce high-quality segmentation across scales.
These results highlight the superiority of PartSAM over existing methods: By replacing heuristic clustering with an automatic promptable framework and leveraging large-scale 3D training, it establishes a more controllable and generalizable model with 3D-aware part understanding.
@@ 5 Conclusion and Discussions
Discussions. Please refer to Appendix Section A.2 for comprehensive ablation studies and additional discussions, including applications, complexity analysis, and limitations.
@@ Appendix A Appendix
A.1 Implementation Details
@@
Model. In all experiments, we follow the setup of Liu et al. (2025) and uniformly sample the input point cloud to 100,000100{,}000 points. For shapes without texture, we assign all points a default gray color. Our encoder is configured with a triplane resolution of $512$, a feature dimension of $448$, and $6$ transformer layers per branch. The number of point patches $N_{c}$ is set to $2,000$. The mask decoder consists of four two-way transformer layers with a feature dimension of $256$ and outputs $3$ masks for a single prompt point.
@@
Training. During training, we simulate the interactive segmentation process as follows: For each ground-truth foreground mask, we first sample the prompt point from the foreground region, selecting points whose distance to the background falls within the bottom third of the total distance range. After obtaining the predicted masks, we iteratively sample subsequent prompt points by selecting the furthest points within the error regions (i.e., the areas where the predicted and ground-truth masks differ). This process is repeated 9 times per mask. The parameter $\lambda$ to balance the triplet loss is set to $0.5$. A comprehensive data augmentation strategy is employed to improve model robustness by introducing controlled variations in geometry, appearance, and scale. The augmentation pipeline includes random rotations, scaling, flipping, and center shifting of the point cloud, as well as chromatic transformations such as auto contrast, translation, and jitter. Overall, PartSAM is trained using the AdamW optimizer for $250k$ iterations with a batch size of $4$ on $8$ NVIDIA H20 GPUs. The learning rate is initialized at $5\times 10^{-4}$ and decayed by a factor of $0.7$ every $50k$ iterations.
@@ A.2 Ablation Study and Discussions
Encoder Variants
w/ Encoder of Point-SAM 45.1 69.8 76.6 79.7 56.4
@@
A.2.1 Ablation of Encoder
To assess the impact of the encoder, we compare three variants of PartSAM trained under identical settings. We begin with the point cloud-based transformer from Point-SAM (Zhou et al., 2025), where point clouds are tokenized and processed by a ViT-initialized transformer (Dosovitskiy et al., 2021). The substantial performance drop on both interactive and automatic tasks underscores the superior scalability of the triplane representation over point-based transformers when handling large-scale 3D data. Next, we employ the pre-trained PartField as our encoder while keeping it frozen. The suboptimal performance indicates that training only the decoder is insufficient for scaling to our dataset, highlighting the critical role of the encoder in building a promptable segmentation model. Finally, we unlock PartField and observe notable improvements, but its performance remains below that of our dual-branch encoder. This gap stems from catastrophic forgetting: direct fine-tuning erodes the strong priors inherited from SAM, reducing segmentation effectiveness. In contrast, our dual-branch design preserves these priors while learning fine-grained semantics from large-scale 3D data, and can further integrate auxiliary inputs (e.g., normals and RGB) to enhance precision.

### v3

@@ Abstract
Segmenting 3D objects into parts is a long-standing challenge in computer vision. To overcome taxonomy constraints and generalize to unseen 3D objects, recent works turn to open-world part segmentation. These approaches typically transfer supervision from 2D foundation models, such as SAM, by lifting multi-view masks into 3D. However, this indirect paradigm fails to capture intrinsic geometry, leading to surface-only understanding, uncontrolled decomposition, and limited generalization. We present PartSAM, the first promptable part segmentation model trained natively on large-scale 3D data. Following the design philosophy of SAM, PartSAM employs an encoder–decoder architecture in which a triplane-based dual-branch encoder produces spatially structured tokens for scalable part-aware representation learning. To enable large-scale supervision, we further introduce a model-in-the-loop annotation pipeline that curates over five million 3D shape–part pairs from online assets, providing diverse and fine-grained labels. This combination of scalable architecture and diverse 3D data yields emergent open-world capabilities: with a single prompt, PartSAM achieves highly accurate part identification, and in a “Segment-Every-Part” mode, it automatically decomposes shapes into both surface and internal structures. Extensive experiments show that PartSAM outperforms state-of-the-art methods by large margins across multiple benchmarks, marking a decisive step toward foundation models for 3D part understanding. Project page: https://czvvd.github.io/PartSAMPage/.
@@ 1 Introduction
Second, we propose an effective encoder that can scale to large 3D datasets. Specifically, we encode shapes into a triplane-based feature field (Chan et al., 2022) using a dual-branch transformer (Vaswani et al., 2017). To better leverage existing priors, we leverage one learnable branch scale to large 3D datasets, while the other frozen path preserves 2D priors of SAM learned through contrastive learning (Liu et al., 2025). Additionally, by incorporating input attributes beyond just coordinates, the model enhances the representation of local shape details.
@@ 2.1 Closed-world 3D Part Segmentation
Early learning-based methods (Qi et al., 2017; Thomas et al., 2019; Wang et al., 2019b; Zhang and Wonka, 2021) typically formulate part segmentation as point-level semantic or instance classification on 3D shapes. These approaches are commonly trained on ShapeNet-Part (Yi et al., 2016) and PartNet (Mo et al., 2019). The limited diversity of these datasets in terms of object and part categories, however, constrains their generalization to unseen data.
@@ 3 Method
An overview of PartSAM is shown in Figure 3. Our goal is to develop a scalable framework for segmenting arbitrary 3D parts, analogous to SAM (Kirillov et al., 2023) in the image domain. The architecture consists of two main components: an input encoder that transforms 3D shapes into structured feature embeddings, and a prompt-guided mask decoder that predicts segmentation masks conditioned on user prompts. We primarily consider 3D click points as prompts, including both positive and negative ones. Formally, given a 3D shape represented as a point cloud $P_{in}\subseteq\mathbb{R}^{N\times d_{in}}$, where each point may include 3D coordinates, surface normals, and optional RGB color (i.e., $d_{in}=9$), and a set of user prompts $P_{prompt}\subseteq\mathbb{R}^{N_{p}\times 3}$, PartSAM first transform the input shape into feature embeddings $F_{c}$ through the input encoder. These are combined with prompt embeddings $F_{p}$ in the mask decoder to generate a binary segmentation mask $M_{out}$. To fully realize this design at scale, we further introduce a data curation pipeline that mines diverse and fine-grained 3D part annotations from large-scale assets, enabling PartSAM to generalize across open-world scenarios. Additional details are provided in Appendix A.1.
@@ 3.2 Prompt-Guided Mask Decoder
For mask decoding, two special tokens are introduced: an output token $T_{out}$ that generates segmentation masks, and an IoU token $T_{iou}$ that estimates mask quality. These tokens, together with the prompt embeddings $F_{p}$, are concatenated and attend bidirectionally to the patch embeddings $F_{c}$ using a two-way transformer (Kirillov et al., 2023):
@@ 3.3 Training Strategy
Following SAM (Kirillov et al., 2023), during training, we simulate interactive segmentation by first sampling a prompt from the center of the ground-truth mask and then iteratively sampling subsequent prompts from prediction error regions for 9 iterations. The IoU prediction is supervised with the MSE loss. The segmentation masks are supervised with a combination of focal loss (Lin et al., 2017) and dice loss (Milletari et al., 2016). For each ground-truth mask, the triplet contrastive loss in Liu et al. (2025) is also used to strengthen the representation ability of the encoder.
$\mathcal{L}=\mathcal{L}_{focal}(M_{out},M_{gt})+\alpha\mathcal{L}_{dice}(M_{out},M_{gt})+\mathcal{L}_{IoU}+\lambda\,\mathcal{L}_{triplet},$ (3)
where $M_{{gt}}$ is the ground-truth mask and $\alpha$ and $\lambda$ are weighting coefficients for the loss terms.
@@ 4.2 Comparison of Class-Agnostic Part Segmentation
As introduced in Section 3.4, PartSAM supports automatic shape decomposition through its “segment every part” mode. We compare PartSAM against five state-of-the-art methods (Liu et al., 2023; Ma et al., 2024; Yang et al., 2024b; Tang et al., 2024; Liu et al., 2025) on this task. We use the PartObjaverse-Tiny (Yang et al., 2024b) and PartNet-E (Liu et al., 2023) datasets for quantitative evaluation and additionally use AI-generated meshes produced by Hunyuan3D (Lai et al., 2025) for qualitative comparison. We follow the experimental protocol from Liu et al. (2025) with a key modification to prevent potential label leakage from mesh connectivity. Specifically, since artist-created meshes in PartObjaverse-Tiny contain strong connectivity cues that align with ground-truth part labels (see Appendix A.2.5), we disable this connectivity information when evaluating PartField (Liu et al., 2025) and instead apply K-Means clustering directly on mesh faces. In contrast, SAMesh (Tang et al., 2024), which fundamentally depends on mesh connectivity through community detection, is evaluated in its original setting. Notably, for baselines that produce multi-scale part segmentations, we evaluate performance by computing the IoU for all candidate masks and reporting the best score. For our PartSAM, multi-scale segmentations are obtained by varying the NMS threshold value $T$ at $0.1$, $0.3$, $0.5$, and $0.7$, respectively.
@@
Qualitative comparisons in Figure 7 further reveal that PartSAM produces robust part segmentations for diverse 3D shapes. SAMesh also exhibits clear boundaries but often generates over-segmented outputs, as its direct projection and merging of 2D SAM masks into 3D space fails to capture underlying geometric structures. SAMPart3D and PartField, which rely on feature clustering, encounter difficulties on complex and detailed structures. Their clustering process frequently yields fragmented or semantically meaningless parts, for example, incorrectly partitioning the body of a grenade into several disjoint segments.
[…]
Figure 9: Qualitative comparison of class-agnostic part segmentation with baselines (Liu et al., 2025; Tang et al., 2024) on AI-generated 3D models (Lai et al., 2025) with interior structures and occluded parts. Each segmented part is represented as a distinct color.
We also conduct a comparison with PartField in the context of hierarchical multi-scale segmentation on AI-generated meshes (Figure 8). PartSAM delivers finer-grained results with sharper boundaries, highlighting its capability to produce high-quality segmentation across scales.
In Figure 9, we further compare PartSAM with baselines on diverse AI-generated meshes with interior or occluded parts. SAMesh benefits from inheriting 2D SAM’s billion-scale mask priors, enabling it to capture extremely fine surface details and produce sharp boundaries (e.g., the car exterior). However, this 2D-driven granularity may also over-fragment large but semantically coherent regions (handbag), and, more importantly, it struggles when parts are occluded in most rendered views—for example, SAMesh cannot reliably recover the objects inside the handbag, the seats and steering wheel inside the car, or the structures hidden beneath the robot’s cloak. In contrast, PartSAM operates directly on native 3D geometry and learns multiscale, semantically meaningful decompositions, producing coherent part groupings for both visible and partially hidden structures.
Overall, these results highlight the superiority of PartSAM over existing methods: By replacing heuristic clustering with an automatic promptable framework and leveraging large-scale 3D training, it establishes a more controllable and generalizable model with 3D-aware part understanding.
@@ 5 Conclusion and Discussions
Discussions. Please refer to the Appendix for comprehensive ablation studies and additional discussions, including applications, complexity analysis, and limitations.
6 Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. T2322012, No. 62572240). It was also supported by the donation from Kerry Group Limited in “30-for-30” Talent Acquisition Campaign of HKUST.
@@ Appendix A Appendix
A.1 Additional Details
A.1.1 Details of the segmentation pipeline
Overall. Given an input point cloud $P_{\text{in}}\in\mathbb{R}^{N\times d}$ and a set of prompt points $P_{\text{prompt}}\in\mathbb{R}^{N_{p}\times 3}$, PartSAM predicts a binary part mask $M_{\text{out}}$.
Encoder. The encoder contains two triplane branches. Each of them follows a similar architecture to Liu et al. (2025). The learnable branch differs only in its input layer, where six additional feature channels (representing XYZ coordinates and normals) are concatenated to the coordinates. Within each branch, a PVCNN (Liu et al., 2019) operating at a voxel resolution of $32^{3}$ first extracts per-point features, which are orthogonally projected onto three axis-aligned planes to form the initial triplane field $F_{\text{plane}}\in\mathbb{R}^{3\times H\times W\times C^{\text{init}}}$. The planes are then downsampled by a factor of $r$ via a two-layer CNN and reshaped into a sequence of size $(3HW/r^{2})\times C^{\text{trans}}$, which is subsequently processed by a transformer. The transformer outputs are upsampled through a transposed convolution layer, reshaped back into three planes, and summed, yielding a triplane representation of size $3\times H\times W\times C$. Point-wise features are sampled from the resulting planes according to their coordinates. A sampling-and-grouping operation (Qi et al., 2017) further aggregates these features to produce the input embeddings $F_{c}\in\mathbb{R}^{N_{c}\times C}$.
Decoder. For a single prompt, we use three output tokens and one IoU token, resulting in three predicted masks. For multiple prompts, we follow SAM (Kirillov et al., 2023) by adding an additional output token that predicts a single aggregated mask, which is ignored in the single-prompt case and used exclusively in the multi-prompt setting. Thus, in both settings, the decoder always receives five special tokens (three output tokens, one IoU token, and one auxiliary output token). Prompt points are encoded by sampling the triplane feature field at their coordinates and adding positional and prompt-type (positive or negative) embeddings, producing $F_{p}\in\mathbb{R}^{N_{p}\times C^{\prime}}$. We project the input embeddings $F_{c}$ to the same dimension $C^{\prime}$ via a linear layer, yielding $F_{c}\in\mathbb{R}^{N_{c}\times C^{\prime}}$. Output tokens and the IoU token are concatenated with $F_{p}$ to form a unified token set with shape of $(N_{p}+5)\times C^{\prime}$, which is processed jointly with $F_{c}$. The two sets interact through a four-layer two-way transformer (similar to Kirillov et al. (2023)), where each layer applies self-attention within tokens, followed by cross-attention from tokens to embeddings and another cross-attention from embeddings back to tokens. This produces refined output token $T_{\text{out}}^{\prime}$, refined IoU token $T_{\text{iou}}^{\prime}$, and updated input embeddings $F_{c}^{\prime}$. The refined input embeddings $F_{c}^{\prime}$ are upsampled to per-point resolution via distance-based interpolation followed by an MLP, yielding a tensor of shape $N\times C^{\prime}$. Then, each refined $C^{\prime}$-dimensional output token predicts mask logits through a dot product with the upsampled embeddings, which is passed through a sigmoid and then thresholded to obtain a binary mask.
IoU head. When a single prompt is given, the refined IoU token $T_{\text{iou}}^{\prime}\in\mathbb{R}^{C}$ is passed through a three-layer MLP to predict mask-wise IoU scores. The mask with the highest score is then selected as the final output $M_{\text{out}}$.
@@ A.1.1 Details of the segmentation pipeline
Hyperparameters. In all experiments, we follow the setup of Liu et al. (2025) and uniformly sample the input point cloud to $N_{p}=100{,}000$ points. For shapes without texture, we assign all points a default gray color. Our encoder is configured with a triplane resolution of $H=W=512$, $C^{init}=64$, $C^{trans}=1024$, $C=448$, and $6$ transformer layers per branch. The number of point patches $N_{c}$ is set to $2,000$. The feature dimension of two-way transformer layers $C^{\prime}$ is set to $256$.
A.1.2 Mechanism of incorporating previous point prompt
In the first round of interaction, we input the first prompt point and obtain the initial mask result. In subsequent rounds, we concat the last predicted mask logits with coordinates and adopt the same sampling-and-grouping strategy as in Equation (1) to extract dense mask embeddings $F_{p}\subseteq\mathbb{R}^{N_{c}\times C}$. These dense embeddings are directly added to the input embeddings $F_{c}$, allowing the model to incorporate information from previous prompts. For the new prompt point, we distinguish positive and negative prompts by adding different learnable prompt-type embeddings to $F_{p}$.
A.1.3 Data Curation Details
Our data curation follows a fully automated two-stage pipeline designed to construct large-scale, high-quality 3D part annotations. In the first stage, several safeguards are applied to filter out low-quality data from Objaverse(-XL). We remove scanned objects that generally exhibit poor geometric fidelity and are unsuitable for part annotation. We further discard parts that are excessively small or large based on their foreground point statistics, and exclude objects containing fewer than 3 or more than 50 parts to maintain reasonable segmentation granularity. In the second stage, we focus exclusively on high-poly, artist-designed meshes that provide rich geometric structure and numerous connected components. Artist-defined connectivity information is incorporated into PartField’s clustering process to produce clean, well-separated segments. A model-in-the-loop procedure then imposes two additional safeguards: (i) a part-level filter that accepts only segments meeting the interactive IoU criterion, and (ii) a shape-level filter that retains only shapes with more than five valid parts. Only shape–part pairs satisfying all conditions are preserved, ensuring that clustered masks lacking clear semantic meaning (e.g., wheels merged into the car body) are consistently removed.
To ensure that the pipeline scales to millions of shape–part pairs, the whole data curation process is designed to be fully automated. Although a very small number of imperfect cases may remain due to the absence of human verification, their proportion is negligible and does not affect model behavior. As evidenced by the scaling curves in Figure 11, model performance continues to improve as dataset size increases, demonstrating both the scalability of the model and the overall high quality of the curated training data.
A.1.4 Training Details
@@
During training, we simulate the interactive segmentation process as follows: For each ground-truth foreground mask, we first sample the prompt point from the foreground region, selecting points whose distance to the background falls within the bottom third of the total distance range. After obtaining the predicted masks, we iteratively sample subsequent prompt points by selecting the furthest points within the error regions (i.e., the areas where the predicted and ground-truth masks differ). This process is repeated 9 times per mask. We set $\alpha=2$ and $\lambda=0.5$. A comprehensive data augmentation strategy is employed to improve model robustness by introducing controlled variations in geometry, appearance, and scale. The augmentation pipeline includes random rotations, scaling, flipping, and center shifting of the point cloud, as well as chromatic transformations such as auto contrast, translation, and jitter. Overall, PartSAM is trained using the AdamW optimizer for $250k$ iterations with a batch size of $4$ on $8$ NVIDIA H20 GPUs. The learning rate is initialized at $5\times 10^{-4}$ and decayed by a factor of $0.7$ every $50k$ iterations.
@@ A.2 Ablation Study and Discussions
Model Variants
Point-SAM 38.9 62.9 71.5 76.8 48.6
w/o pre-trained weights 48.3 73.6 79.1 80.5 60.5
@@
A.2.1 Ablation of Model Architecture
To assess the impact of the model architecture, we compare three variants of PartSAM trained under identical settings.
Comparison with Point-SAM. To comprehensively compare with the existing promptable segmentation model Point-SAM (Zhou et al., 2025), we train it using our curated dataset and report the results in Table 3. The significant performance drops on both the interactive and automatic segmentation tasks indicate that Point-SAM fails to scale on our native 3D data. Compared with Point-SAM, our method offers advantages in both network architecture and training scheme for the part segmentation task. Regarding architecture, Point-SAM uses a transformer encoder operating on unstructured point clouds, whereas our method performs transformer modeling on a more advanced triplane feature field. The dual-branch encoder also brings in 2D SAM priors, which improve generalization and stabilize training on large-scale 3D data. We also benefit from the training scheme: the triplet contrastive loss strengthens part-aware feature learning, and its formulation is naturally robust to the scale and granularity ambiguity of 3D parts. Together, the continuous feature representation and the contrastive learning scheme give our framework substantially stronger scalability.
Analysis of Pre-trained Encoder. Then, we evaluate a variant that trains PartSAM entirely from scratch, without using any initialization. This variant performs worse in terms of all tasks, indicating that the SAM-derived priors provide useful part-aware cues while not introducing lifting-related artifacts. Next, we employ the pre-trained PartField as the encoder while keeping it frozen. The suboptimal performance under this setting shows that training only the decoder is insufficient for scaling to our dataset, highlighting the importance of the encoder in building a promptable segmentation model. We then fine-tune the entire PartField encoder and observe notable improvements; however, its performance remains below that of our dual-branch encoder. This gap mainly stems from catastrophic forgetting—direct fine-tuning erodes the strong priors inherited from SAM, leading to degraded segmentation quality. In contrast, our dual-branch design preserves these 2D SAM priors in a frozen branch, while the learnable 3D-native branch acquires fine-grained 3D semantics from large-scale training. This architecture avoids forgetting effects and supports integrating auxiliary inputs (e.g., normals or RGB), resulting in the best overall performance.
Figure 10: Visual comparison of native 3D part labels with 2D lifting part labels.